OpenStudy (thefluffmuncher):

7. What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8 (Points : 4) parallel perpendicular neither parallel nor perpendicular they are the same line

OpenStudy (thefluffmuncher):

not sure but i think it's c

OpenStudy (dtan5457):

Simplify the first one to 6y=8-3x y=8/6-1/2x y=2x-8 First one's slope is -1/2x and 2nd slope is 2x Parallel=same slope perpendicular=reciprocal and negation of original slope Which one is it?

OpenStudy (dtan5457):

"perpendicular=reciprocal and negation of original slope" For example, if a slope was 1/5, the perpendicular slope would be -5/1

OpenStudy (thefluffmuncher):

so it's perpendicular

OpenStudy (dtan5457):

yes.

OpenStudy (thefluffmuncher):

8. Which of the following has a slope of –2 and a y-intercept of 4? (Points : 4) y = –2x + 4 y = 2x + 4 y = 2x – 4 y = –2x – 4

OpenStudy (thefluffmuncher):

i'm pretty sure it's a on this one

OpenStudy (dtan5457):

Correct.

OpenStudy (dtan5457):

Anymore questions?

OpenStudy (thefluffmuncher):

don't think so thanks for the useful help :)

OpenStudy (dtan5457):

Np. Feel free to medal and fan and tag me into future questions :)